Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

A card shop contains 5 birthday cards, 5 holiday cards...

Expert replies
by BTGmoderatorLU » Mon Mar 19, 2018 2:43 pm
A card shop contains 5 birthday cards, 5 holiday cards, and 5 graduation cards. If three cards are purchased at random from the shop, what is the probability that the three cards will be of the same type?

A. 6/91
B. 5/93
C. 4/95
D. 3/97
E. 2/99

The OA is A.

I'm confused by this PS question. Experts, any suggestion about how can I solve it? Thanks in advance.
Join the discussion
Source: — Problem Solving |

by [email protected] » Mon Mar 19, 2018 2:54 pm
Hi LUANDATO,

We're told that a card shop contains 5 birthday cards, 5 holiday cards, and 5 graduation cards and that three cards are purchased at random from the shop. We're asked for the probability that the three cards will be of the SAME type.

To start, the first card can be ANY of the 15 cards - it's the second and third cards that will have to match the first.

Probability that the second card matches the first = (4/14)
Probability that the third card matches the first = (3/13)

(4/14)(3/13) = 12/182 = 6/91

Final Answer: A

GMAT assassins aren't born, they're made,
Rich
Contact Rich at [email protected]
Image
Join the discussion

by Jeff@TargetTestPrep » Wed Mar 21, 2018 3:23 pm
LUANDATO wrote:A card shop contains 5 birthday cards, 5 holiday cards, and 5 graduation cards. If three cards are purchased at random from the shop, what is the probability that the three cards will be of the same type?

A. 6/91
B. 5/93
C. 4/95
D. 3/97
E. 2/99
The number of ways to select 3 birthday cards is 5C3 = (5 x 4 x 3)/(3 x 2) = 10, which is also the number of ways to select 3 holiday cards and the number of ways to select 3 graduation cards.

The total number ways to select 3 cards from 15 is:

15C3 = (15 x 14 x 13)/(3 x 2) = 5 x 7 x 13 = 455

So the probability of selecting 3 birthday cards or 3 holiday cards or 3 graduation cards is:

10/455 + 10/455 + 10/455 = 30/455 = 6/91

Answer: A

Jeffrey Miller
Head of GMAT Instruction
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews
Join the discussion

by BTGmoderatorRO » Sun Mar 25, 2018 6:25 am
There are 5 Birthday cards (B), 5 Holiday cards (H) and 5 graduation cards (G).
Total number of cards=15
prob. of purchasing three cards of the same type can be pr(BBB), pr(HHH) or pr(GGG).
Therefore,
$$pr\left(BBB\right)+pr\left(HHH\right)+pr\left(GGG\right)$$
$$pr\left(\frac{5}{15}\cdot\frac{4}{14}\cdot\frac{3}{13}\right)+pr\left(\frac{5}{15}\cdot\frac{4}{14}\cdot\frac{3}{13}\right)+pr\left(\frac{5}{15}\cdot\frac{4}{14}\cdot\frac{3}{13}\right)$$
$$pr\left(\frac{1}{3}\cdot\frac{2}{7}\cdot\frac{3}{13}\right)+pr\left(\frac{1}{3}\cdot\frac{2}{7}\cdot\frac{3}{13}\right)+pr\left(\frac{1}{3}\cdot\frac{2}{7}\cdot\frac{3}{13}\right)$$
$$\frac{6}{273}+\frac{6}{273}+\frac{6}{273}=\frac{18}{273}=\frac{6}{91}$$
Hence, option A is the correct answer
Join the discussion

by Brent@GMATPrepNow » Sun Mar 25, 2018 6:54 am
LUANDATO wrote:A card shop contains 5 birthday cards, 5 holiday cards, and 5 graduation cards. If three cards are purchased at random from the shop, what is the probability that the three cards will be of the same type?

A. 6/91
B. 5/93
C. 4/95
D. 3/97
E. 2/99
Let's first rewrite our probability and then apply probability rules.

P(All 3 cards the same type) = P(1st card is ANY type AND 2nd card matches type of 1st card AND 3rd card matches type of 1st card)
= P(1st card is ANY type) x P(2nd card matches type of 1st card) x P(3rd card matches type of 1st card)
= 1 x 4/14 x 3/13
= 6/91

Answer: A

ASIDE
P(1st card is ANY type) =1 because the first selection can be any type
P(2nd card matches type of 1st card) = 4/14, because once the 1st card is selected, there are 14 cards remaining, and there are 4 cards left that are the same type as the first card
P(3rd card matches type of 1st card) = 3/13, because once cards 1 and 2 have been selected, there are 13 cards remaining, and only 3 of them are the same type as the first card

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion